OpenStudy (mergirlblue):

Please help ASAP! Let f(x)=x^2+5 and g(x)=x+5/x . find (g ∘ f) (-3). Thanks!

RhondaSommer (rhondasommer):

can help

OpenStudy (mergirlblue):

cool :)

RhondaSommer (rhondasommer):

one sec tho i need to restart my comp. its being glitchy

OpenStudy (samigupta8):

At first u need to find value of f(-3) which will cum out to be (-3)^2+5 which is 14

OpenStudy (samigupta8):

Then u can similarly find the value for g(14l

OpenStudy (samigupta8):

Sorry it was meant to be g(14)

OpenStudy (mergirlblue):

My answer choices are 1/4, 14/19, 19/14, 49/9 Is the answer 14/19 ?

OpenStudy (michele_laino):

hint: we can write, by definition of compposed function: \[\large \left( {g \circ f} \right)\left( 3 \right) = g\left( {f\left( { - 3} \right)} \right) = f\left( { - 3} \right) + \frac{5}{{f\left( { - 3} \right)}} = ...?\]

OpenStudy (mergirlblue):

so it's 19/14 ?

OpenStudy (michele_laino):

sorry I have made a typo, here is the right formula: \[\Large \begin{gathered} \left( {g \circ f} \right)\left( 3 \right) = g\left( {f\left( { - 3} \right)} \right) = \frac{{f\left( { - 3} \right) + 5}}{{f\left( { - 3} \right)}} = \hfill \\ \hfill \\ = \frac{{14 + 5}}{{14}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, you are right!

OpenStudy (mergirlblue):

Yay! Thank you

OpenStudy (michele_laino):

:)

OpenStudy (mergirlblue):

Thank you everybody